# Privacy Reasoning in Ambiguous Contexts

### v1

@@ 1 Introduction
Personal AI agents are expected to interact with a user to complete tasks on their behalf. To effectively complete a broad variety of tasks in a personalized manner, such agents require access to significant amounts of personal and potentially sensitive user data. This implies that a personal agent must be able to safeguard this personal information while achieving high utility. To be truly useful, the agent must therefore be able to judge based on the context of the interaction when it is appropriate (as defined in the contextual integrity framework nissenbaum2004privacy ; nissenbaum2009privacy ) and consistent with a specific user’s expectations to share personal information in interactions with third-party tools and agents.
@@
Figure 1: Overview of the agent completing personalized tasks on behalf of users. The Camber framework enables context clarification that allows the agent to judge the appropriateness of sharing personal information in various scenarios that are ambiguous and incompletely specified.
In this paper, we dive deeper into the privacy judgments of LLMs from the angle of personal agents, where the user could provide insufficient yet trusted context. Figure 1 highlights this setup, where the agent needs to perform privacy assessments on behalf of users, and these judgments are based on ambiguous contexts. Specifically, we investigate the capacity of LLMs to make data sharing decisions, the impact of situational information completeness—i.e. contextual ambiguity—on the accuracy of these judgments, and the effectiveness of disambiguating context.
@@
To mitigate this effect, we develop Camber, a context disambiguation framework to simulate the effect of clarifications that the agent would receive in practice (e.g., through questions to the users or database queries). Drawing inspiration from the principles of contextual integrity theory, Camber automatically generates plausible expansions of ambiguous scenarios to obtain a more fully specified context. Using Camber, we first identify the sources of ambiguity in the existing scenarios, for which disambiguation is crucial to improve LLM privacy reasoning. We then expand existing contextual facts to cover ambiguities and generate synthetic examples. Our experiments with disambiguated scenarios demonstrate that these are easier for models to reason on, resulting in consistently higher privacy assessment performance and reduced sensitivity to prompt variability. Our study reveals that addressing ambiguous contexts plays a central role in developing practical agents that can be deployed in production.
In summary, the key contributions of this paper are as follows: (1) identifying and studying the concept of contextual ambiguity as a barrier to performance; (2) creating synthetic datasets with ambiguous contexts for measuring prompt sensitivity and consistency; (3) introducing Camber, a disambiguation framework accommodating contextual integrity theory and reasoning-guided approaches; and (4) showing that systematically disambiguating context leads to significant performance improvements and reductions in prompt sensitivity.
@@ 3 Problem Statement
We consider an agent, depicted in Figure 1, modeled by an LLM $\phi$, tasked with making a binary decision about sharing a piece of information (e.g., a medical condition or an appointment time). Such information is served to the model through data structures called scenarios $\mathcal{S_{C}}$, parameterized by a set of contextual fields $\mathcal{C}$, often inspired by the contextual integrity theory nissenbaum2004privacy ; nissenbaum2009privacy . The task of the model $\phi$ is then to output $\phi(\mathcal{S_{C}})\in\{\texttt{yes},\texttt{no}\}$, where yes (respectively no) corresponds to the conclusion that it is appropriate (respectively not) to share the data in the scenario $\mathcal{S_{C}}$. See Figure 2 for an example. The agent can interact with the environment to acquire evidence for determining the appropriateness of sharing.
@@
Context clarification. We focus on studying what type of clarifying context is required to improve privacy assessments. We intentionally allow for a broad notion that covers settings where the agent may ask clarifying questions to the user, or access more fine-grained information about the sharing scenario. We develop a framework to systematically expand $\mathcal{C}$ with various types of new context $\mathcal{N}$, and observing its effect on privacy judgments through $\mathcal{S_{C|N}}$.
Performance metrics. Given a labeled dataset $D=\{\mathcal{(S_{C}}_{i},l_{i})\}_{i}$, with $l_{i}\in\{\texttt{yes},\texttt{no}\}$, we evaluate the model’s performance based on precision (P), recall (R), and $F_{1}$ score. Precision is defined as $P:=\frac{|i:\theta(\mathcal{S_{C}}_{i})=\texttt{yes}\wedge l_{i}=\texttt{yes}|}{|i:\theta(\mathcal{S_{C}}_{i})=\texttt{yes}|}$, recall is defined as $R:=\frac{|i:\theta(\mathcal{S_{C}}_{i})=\texttt{yes}\wedge l_{i}=\texttt{yes}|}{|i:l_{i}=\texttt{yes}|}$, and $F_{1}:=\frac{P\cdot R}{P+R}$ is their harmonic mean. High precision indicates that the model is effective at minimizing false positives, i.e., not sharing data that it shouldn’t share. In contrast, high recall means that the model is good at identifying positive instances, i.e., sharing when appropriate. Therefore, a model that is overly conservative about what data to share will have high precision and low recall, whereas a model that is more permissive will have higher recall.
Threat model. We assume a non-adversarial setting. Neither the prompt nor the provided context $\mathcal{C|N}$ are intentionally malicious nor designed to deceive the model through adversarial manipulation.
4 Contextual Ambiguity Analysis
@@
To better evaluate both the privacy and utility of LLMs in probing-style tasks, we use two existing privacy benchmarks: PrivacyLens salt-nlp2024privacylens and ConfAIde mireshghallah2023confaide . PrivacyLens contains 493 examples characterized by five contextual integrity fields: data type, data subject, data sender, data recipient and transmission principle. In ConfAIde, examples describe scenarios where an aware agent tries to disclose a sensitive personal detail about a subject agent to an oblivious agent for a reveal reason. Each of the 270 ConfAIde examples contains seven fields (the subject agent’s aware-agent-relationship and oblivious-agent-relationship, in addition to the five above), which are expanded into short descriptive stories on which privacy assessments are performed. Because the two datasets contain only inappropriate-to-share scenarios (denoted as $\mathcal{S^{-}}$) to measure privacy leakage, we augment them by incorporating instances where data is appropriate to share ($\mathcal{S^{+}}$). This allows us to perform comprehensive assessments of utility. For each $\mathcal{S^{-}}$, we generate a corresponding positive scenario $\mathcal{S^{+}}$ by modifying a subset of its fields. For PrivacyLens, we source these appropriate-to-share fields directly from the dataset, while for ConfAIde we generate them using an LLM. We validate the positive examples manually. The resulting augmented datasets, termed PrivacyLens+ and ConfAIde+, incorporate $\mathcal{S^{+}}$ alongside the original $\mathcal{S^{-}}$, and contain 986 and 540 examples respectively. A detailed description of the dataset generation procedure is found in Appendix D.1 and  D.2.
@@
We use three state-of-the-art models: two reasoning models (Gemini 2.5 Pro, version Gemini 2.5 Pro Experimental, and Claude 3.7 Sonnet, version claude-3-7-sonnet-20250219), and one flagship non-reasoning model (GPT-4.1, version gpt-4.1-2025-04-14) [^1] We set temperature to 0 when evaluating all models, a setting unavailable in OpenAI’s reasoning o3 model.. Our goal is to investigate privacy assessments across multiple frontier models, rather than to compare their performance.
@@
Following existing benchmarks mireshghallah2023confaide , we start with a neutral prompt that asks “is it okay” to share the relevant data without providing more information or context. Next, we create a restrictive prompt variant by adding “Please answer the question considering privacy norms in the response” to the neutral prompt, akin to prior work mireshghallah2023confaide . Finally, we create a permissive prompt variant by adding “Please answer the question sharing as much as possible while respecting privacy.” The goal of these variants is to identify how performance changes across distinct input instructions, rather than to develop the best prompt variants for each setting. The exact prompts we used for PrivacyLens+ and ConfAIde+ are in Appendix E. For all experiments, we set temperature to 00 and the maximum output length to one token (the yes/no judgment).
Results.
Table 1 highlights the performance for each model on the three prompt variants across datasets. In general, the neutral prompt yields high precision (85.6–92.2% for PrivacyLens+ and 98.6–100.0% for ConfAIde+) but lower recall (62.7–75.7% for PrivacyLens+ and 85.6–95.6% for ConfAIde+); i.e., the models preserve privacy but have lower utility. High precision is consistent with prior work, which observed that negative examples are more easily distinguishable by recent models sun2024trustllm , but our experiments show that this comes at the cost of low recall.
Both the restrictive and permissive prompt variants increase precision, but this generally comes at a significant cost in terms of recall. In the most extreme case, switching to the restrictive prompt variant with Claude on the ConfAIde+ dataset increases precision from 99.0% to 100% but drops recall from 85.5% to 49.6%. This utility drop is partially recovered with the permissive prompt, but typically does not reach the performance of the neutral prompt; i.e., the $F_{1}$ score remains lower. Overall, these results indicate high prompt sensitivity across models. Furthermore, in only two cases does the permissive prompt increase recall over the neutral one (Claude on PrivacyLens+ and GPT-4.1 on ConfAIde+). This shows that it is challenging to develop prompt variants that bias the models to achieve high privacy and utility simultaneously, as they tend to be overly conservative in prioritizing privacy over utility.
Model Intent of ConfAIde+ PrivacyLens+
@@ Results.
Gemini 2.5 Pro neutral $100.0$ $73.7$ $84.9$ $86.5$ $69.0$ $76.8$
restrictive $100.0$ $57.0$ $72.6$ $91.3$ $40.6$ $56.2$
permissive $100.0$ $68.9$ $81.6$ $88.9$ $63.5$ $74.1$
GPT-4.1 neutral $99.0$ $75.9$ $86.0$ $85.6$ $75.7$ $80.3$
restrictive $99.5$ $68.9$ $81.4$ $93.2$ $61.5$ $74.1$
permissive $99.5$ $78.5$ $87.8$ $88.3$ $73.4$ $80.2$
Claude 3.7 Sonnet neutral $99.0$ $75.2$ $85.5$ $92.2$ $62.7$ $74.6$
restrictive $100.0$ $33.0$ $49.6$ $95.6$ $43.8$ $60.1$
permissive $98.6$ $53.7$ $69.5$ $92.7$ $66.5$ $77.4$
@@
Table 1: The precision (P), recall (R), and $F_{1}$ score of each model in classifying the appropriateness of information flows across the ConfAIde+ and PrivacyLens+ datasets, using three different prompt variants. Our names for the prompt variants capture the intent behind them, which we can see is not always the same as the effect they produced.
Analysis.
@@
These results and the underlying high prompt sensitivity show that it remains challenging to tune prompts to achieve high classification accuracy, considering both privacy and utility. An inspection of the underlying data reveals that many of the misclassified examples appear underspecified and ambiguous, meaning that the described situation allows both an interpretation where sharing is appropriate and where it is not. Consider the PrivacyLens example in Figure 2 that is labeled as inappropriate to share. The rationale provided by the model reflects the fact that there are several plausible interpretations of what insurance information refers to, whether the coworker is authorized or has a need to know, and whether the communication channel is approved for such information, all of which affect the appropriateness of sharing. Appendix A lists additional examples with ambiguous context.
@@
While systematic prompt tuning can likely improve overall performance, this analysis also suggests contextual ambiguity as a key barrier. However, the methods for reducing context ambiguity and their resulting impact remain unclear. We investigate this issue in the following section.
5 Camber: the Context Disambiguation Framework
@@
The contexts in which LLM-based agents are required to make privacy assessments are inevitably vague. In the previous section, we saw that models perform poorly in the face of contextual ambiguity, in terms of being overly conservative and exhibiting high prompt sensitivity. We therefore consider a model where agents have the opportunity to clarify this context (e.g., by asking a user or querying a database). To understand the gains that context clarification could bring, this section introduces the Camber  [^2] We name the framework Camber, for it reflects how in engineering, a subtle camber has a significant performance impact, mirroring how the framework’s targeted clarification of contexts yields substantial accuracy gains. disambiguation framework, and the three synthetic context disambiguation strategies implemented using Camber, as shown in Figure 3.
The first strategy corresponds to an agent that obtains arbitrary additional information contextually relevant to the privacy assessment (e.g., clarifying a request for insurance information with details that the request is about the policy numbers and limits). We call this the label-independent expansion strategy because the new context does not directly relate to the appropriateness label. In the second strategy, the agent obtains evidence directly indicative of the appropriateness label (e.g., retrieving an email that states the insurance information request also includes social security numbers). We call this the label-dependent expansion strategy because the new context biases the model with appropriateness information. Finally, the agent might reason about which fields in the information request are ambiguous and seek evidence to reduce this ambiguity (e.g., reasoning that sharing insurance information requires the consent of policyholders, and asking the user whether consent was provided). We call this the reasoning-guided expansion strategy (Section 5.2). To evaluate these strategies and their impact on improving privacy assessments, in Section 5.1 we describe the mechanisms enabling Camber to generalize across both the ConfAIde+ and PrivacyLens+ datasets. A more extended description of the Camber context disambiguation framework can be found in Appendix C.

### v2

@@ 1 Introduction
Personal AI agents are expected to interact with a user to complete tasks on their behalf. To effectively complete a broad variety of tasks in a personalized manner, such agents require access to significant amounts of personal and potentially sensitive user data. This implies that an agent must be able to safeguard this personal information while achieving high utility. To be truly useful, the agent must therefore be able to judge based on the context of the interaction when it is appropriate (as defined in the contextual integrity framework nissenbaum2004privacy; nissenbaum2009privacy) and consistent with a specific user’s expectations to share personal information in interactions with third-party tools and agents.
@@
Figure 1: Left: The Camber framework studies the impact of ambiguity on privacy judgments and highlights the benefits of resolving it through various disambiguation strategies. Right: Camber aims to inform the design of context clarification mechanisms for future agents completing personalized privacy tasks on behalf of users.
In this paper, we dive deeper into the privacy judgments of LLMs from the angle of personal agents, where the user could provide insufficient yet trusted context. Figure 1 highlights this setup, where the agent needs to perform privacy judgments on behalf of users, and these judgments are based on ambiguous contexts. Specifically, we investigate the capacity of LLMs to make data sharing decisions, the impact of situational information completeness—i.e. context ambiguity—on the accuracy of these judgments, and the effectiveness of disambiguating context.
@@
To mitigate this effect, we develop Camber, a context disambiguation framework to simulate the effect of clarifications that the agent would receive in practice (e.g., through questions to the users or database queries). Drawing inspiration from the principles of contextual integrity theory, Camber automatically generates plausible expansions of ambiguous scenarios to obtain a more fully specified context. Using Camber, we first identify the sources of ambiguity in the existing scenarios, for which disambiguation is crucial to improve LLM privacy reasoning. We then expand existing contextual facts to cover ambiguities and generate synthetic examples. Our experiments with disambiguated scenarios demonstrate that these are easier for models to reason on, resulting in consistently higher privacy judgments performance and reduced sensitivity to prompt variability. Our study reveals that addressing ambiguous contexts plays a central role in developing practical agents that can be deployed in production.
In summary, the key contributions of this paper are as follows: (1) identifying and studying the concept of context ambiguity as a barrier to performance; (2) creating synthetic datasets with ambiguous contexts for measuring prompt sensitivity and consistency; (3) introducing Camber, a disambiguation simulation framework accommodating contextual integrity theory and reasoning-guided approaches; and (4) showing that systematically disambiguating context leads to significant performance improvements and reductions in prompt sensitivity.
@@ 3 Problem Statement
We consider an agent, depicted in Figure 1, modeled by an LLM $\phi$, tasked with making a binary judgment about sharing a piece of information (e.g., a medical condition or an appointment time). Such information is served to the model through data structures called scenarios $\mathcal{S_{C}}$, parameterized by a set of contextual fields $\mathcal{C}$, often inspired by the contextual integrity theory nissenbaum2004privacy; nissenbaum2009privacy. The task of the model $\phi$ is then to output $\phi(\mathcal{S_{C}})\in\{\texttt{yes},\texttt{no}\}$, where yes (respectively no) corresponds to the conclusion that it is appropriate (respectively not) to share the data in the scenario $\mathcal{S_{C}}$. See Figure 2 for an example.
@@
Context clarification. In particular, we consider an agent that can seek additional information to clarify the context of a request before rendering a privacy judgment. We intentionally allow for a broad range of clarification options, including settings where the agent may ask clarifying questions to the user, or access more fine-grained environment information about the sharing scenario. To inform the design of such a context clarification process, we focus in this paper on studying what type of clarifying context improves privacy judgments. To this end, we develop a simulation framework to systematically expand $\mathcal{C}$ with various types of new context $\mathcal{N}$, and observe its effect on privacy judgments $\phi(\mathcal{S_{C|N}})$.
Performance metrics. Given a labeled dataset $D=\{\mathcal{(S_{C}}_{i},l_{i})\}_{i}$, with $l_{i}\in\{\texttt{yes},\texttt{no}\}$, we evaluate the model’s performance based on precision (P), recall (R), and $F_{1}$ score. Precision is defined as $P:=\frac{|i:\phi(\mathcal{S_{C}}_{i})=\texttt{yes}\wedge l_{i}=\texttt{yes}|}{|i:\phi(\mathcal{S_{C}}_{i})=\texttt{yes}|}$, recall is defined as $R:=\frac{|i:\phi(\mathcal{S_{C}}_{i})=\texttt{yes}\wedge l_{i}=\texttt{yes}|}{|i:l_{i}=\texttt{yes}|}$, and $F_{1}:=2\cdot\frac{P\cdot R}{P+R}$ is their harmonic mean. High precision indicates that the model is effective at minimizing false positives, i.e., not sharing data that it shouldn’t share. In contrast, high recall means that the model is good at identifying positive instances, i.e., sharing when appropriate. Therefore, a model that is overly conservative about what data to share will have high precision and low recall, whereas a model that is more permissive will have higher recall.
Threat model. We assume a non-adversarial setting. Neither the prompt nor the provided context $\mathcal{C|N}$ are malicious or designed to deceive the model through adversarial manipulation.
4 Context Ambiguity Analysis
@@
To better evaluate both the privacy and utility of LLMs in probing-style tasks, we use two existing privacy benchmarks: PrivacyLens salt-nlp2024privacylens and ConfAIde mireshghallah2023confaide. PrivacyLens contains 493 examples characterized by five contextual integrity fields: data type, data subject, data sender, data recipient and transmission principle. In ConfAIde, examples describe scenarios where an aware agent tries to disclose a sensitive personal detail about a subject agent to an oblivious agent for a reveal reason. Each of the 270 ConfAIde examples contains seven fields (the subject agent’s aware-agent-relationship and oblivious-agent-relationship, in addition to the five above), which are expanded into short descriptive stories on which privacy judgments are performed. Because the two datasets contain only inappropriate-to-share scenarios (denoted as $\mathcal{S^{-}}$) to measure privacy leakage, we augment them by incorporating instances where data is appropriate to share ($\mathcal{S^{+}}$). This allows us to perform comprehensive assessments of utility. For each $\mathcal{S^{-}}$, we generate a corresponding positive scenario $\mathcal{S^{+}}$ by modifying a subset of its fields. For PrivacyLens, we source these appropriate-to-share fields directly from the dataset, while for ConfAIde we generate them using an LLM. We validate the positive examples manually. The resulting augmented datasets, termed PrivacyLens+ and ConfAIde+, incorporate $\mathcal{S^{+}}$ alongside the original $\mathcal{S^{-}}$, and contain 986 and 540 examples respectively. A detailed description of the dataset generation procedure is found in Appendix D.
@@
We use three state-of-the-art models: two reasoning models (Gemini 2.5 Pro, version Gemini 2.5 Pro Experimental, and Claude 3.7 Sonnet, version claude-3-7-sonnet-20250219), and one flagship non-reasoning model (GPT-4.1, version gpt-4.1-2025-04-14). [^1] We set temperature to 0 when evaluating all models, a setting unavailable in OpenAI’s reasoning o3 model. Our goal is to investigate privacy judgments across multiple frontier models, rather than to compare their performance.
@@
Following existing benchmarks mireshghallah2023confaide, we start with a neutral prompt that asks “is it okay” to share the relevant data without providing more information or context. Next, we create a restrictive prompt variant by adding “Please answer the question considering privacy norms in the response” to the neutral prompt, akin to prior work mireshghallah2023confaide. Finally, we create a permissive prompt variant by adding “Please answer the question sharing as much as possible while respecting privacy.” The goal of these variants is to identify how performance changes across distinct input instructions, rather than to develop the best prompt variants for each setting. In addition to these variants that change privacy directives in the prompts, we also evaluate the model’s sensitivity to paraphrasing-style perturbations in Appendix B.2.4. The complete prompts are listed in Appendix E. For all experiments, we set temperature to 0 and the maximum output length to one token (the yes/no judgment).
Model performance is sensitive to prompts.
Table 1 highlights that model performance is highly sensitive to the prompt variant used across datasets. The neutral prompt generally yields high precision (85.6–92.2% for PrivacyLens+ and 98.6–100.0% for ConfAIde+) but lower recall (62.7–75.7% for PrivacyLens+ and 85.6–95.6% for ConfAIde+), indicating models prioritize privacy over utility, achieving high precision at the cost of low recall. The high precision also aligns with prior work finding that models more easily distinguish negative examples sun2024trustllm.
These restrictive and permissive prompt variants exacerbate the trade-off between precision and recall. In the most extreme case, switching to the restrictive prompt for Claude on ConfAIde+ drops recall from 85.5% to 49.6% for a marginal 1% gain in precision. While the permissive prompt sometimes recovers this utility loss, it rarely surpasses the neutral prompt’s performance, resulting in a lower overall $F_{1}$ score. In fact, the permissive prompt only increased recall over the neutral one in two of six cases (Claude on PrivacyLens+ and GPT-4.1 on ConfAIde+). This sensitivity extends beyond the privacy directive to the phrasing itself; as detailed in Appendix B.2.4, simple rephrasing alone caused $F_{1}$ score differences of up to 9.8%. Overall, these results show that it is challenging to craft prompts that achieve both high privacy and high utility, as models tend to be overly conservative in prioritizing privacy over utility.
Model Intent of PrivacyLens+ ConfAIde+
@@
Gemini 2.5 Pro neutral $86.5$ $69.0$ $76.8$ $100.0$ $73.7$ $84.9$
restrictive $91.3$ $40.6$ $56.2$ $100.0$ $57.0$ $72.6$
permissive $88.9$ $63.5$ $74.1$ $100.0$ $68.9$ $81.6$
GPT-4.1 neutral $85.6$ $75.7$ $80.3$ $99.0$ $75.9$ $86.0$
restrictive $93.2$ $61.5$ $74.1$ $99.5$ $68.9$ $81.4$
permissive $88.3$ $73.4$ $80.2$ $99.5$ $78.5$ $87.8$
Claude 3.7 Sonnet neutral $92.2$ $62.7$ $74.6$ $99.0$ $75.2$ $85.5$
restrictive $95.6$ $43.8$ $60.1$ $100.0$ $33.0$ $49.6$
permissive $92.7$ $66.5$ $77.4$ $98.6$ $53.7$ $69.5$
@@
Table 1: The precision (P), recall (R), and $F_{1}$ score of each model in classifying the appropriateness of information flows across the PrivacyLens+ and ConfAIde+ datasets, using three different prompt variants. The names for the prompt variants capture the intent behind them, which we can see is not always the same as the effect they produced.
Context ambiguity contributes to prompt sensitivity.
@@
A manual inspection of the underlying data reveals that many of the misclassified examples appear ambiguous, meaning that the described situation allows both an interpretation where sharing is appropriate and where it is not. To quantify context ambiguity, we additionally conduct an entropy analysis on the model’s privacy judgments sampled 100 times at a higher temperature of 1. Sampled LLM outputs show lower entropy for inappropriate examples compared to the appropriate ones, indicating the model is more consistent when denying data sharing (inappropriate) than when permitting it (appropriate). This is aligned with our findings in Table 1, where models consistently achieve higher precision than recall. Independent manual labeling by the authors confirmed this pattern, showing lower disagreement on inappropriate examples and validating our use of LLM entropy as a proxy for context ambiguity; Section 6 provides a detailed analysis. Analyzing the example in Figure 2 that is labeled as inappropriate to share, we see that the rationale provided by the model reflects several plausible interpretations of what insurance information refers to, whether or not the coworker is authorized or has a need to know, and whether or not the communication channel is approved for such information, all of which affect the appropriateness of sharing. Appendix A lists additional examples with ambiguous context.
@@
While systematic prompt tuning can likely improve overall performance, this analysis also suggests context ambiguity as a key barrier. However, the methods for reducing context ambiguity and their resulting impact remain unclear. We investigate this issue in the following section.
5 Camber: A Context Disambiguation Framework
@@
The contexts in which LLM-based agents are required to make privacy judgments are inevitably vague. In the previous section, we saw that models perform poorly in the face of context ambiguity, in terms of being overly conservative and exhibiting high prompt sensitivity. We therefore consider a model where agents have the opportunity to clarify this context (e.g., by asking a user or querying a database). To understand the gains that context clarification could bring, this section introduces the Camber disambiguation framework, with three synthetic context disambiguation strategies implemented, as shown in Figure 3. These strategies are designed to illuminate whether adding clarifications aids model performance and what type of clarifications are most helpful.
The first strategy serves as a baseline and adds arbitrary information to one of the data fields (e.g., clarifying the data subject of a request from “alumnus” to “an alumnus who had been recognized for contributions in environmental science”), helping us measure to what extent any additional privacy-agnostic information is useful. To ensure that the added information remains in agreement with the existing appropriateness label of the example, we prompt the model to generate only neutral information, in line with prior work salt-nlp2024privacylens. In the second baseline strategy, we aim to synthetically add information that is more privacy-related and indicative of appropriateness to further resolve ambiguity (e.g., retrieving information that the student record was marked as “do not disclose”). To ensure that this added information remains in agreement with the appropriateness label, we include the label in the data generation prompt. We call the first strategy label-independent and the second label-dependent. Finally, our third strategy mimics an agent that reasons about which fields in the information request are ambiguous and seeks evidence to reduce this ambiguity (e.g., reasoning that the sender is not authorized to share alumni information externally). We call this the reasoning-guided expansion strategy (Section 5.2). Comparing results from the first two strategies reveals how important it is for agents to clarify privacy-specific aspects compared to asking general questions about the context. The reasoning-guided strategy further intends to shed light on what types of privacy relevant information would be most effective to acquire. To evaluate these strategies and their impact on improving privacy judgments, in Section 5.1 we describe the mechanisms enabling Camber to generalize across both the ConfAIde+ and PrivacyLens+ datasets. A detailed description of the Camber framework can be found in Appendix C.